If 𝑅 𝑒 = 𝑒 βˆ’ 𝑒 2 𝑖 + 2𝑒 3 𝑗 βˆ’ 3π‘˜, find �𝑑� οΏ½οΏ½ οΏ½οΏ½ Χ¬ .οΏ½οΏ½ 𝑏. ΰΆ± 1 2 𝑅(𝑒)𝑑�


1
Expert's answer
2021-05-11T11:25:22-0400

Ans:-

"𝑅 (𝑢 )= 𝑢 − 𝑢 2 𝑖 + 2𝑢 3 𝑗 − 3𝑘\\"


"\dfrac{dR}{du}=\lim_{u\to0}\dfrac{R(u+\Delta u-R(u))}{\Delta u}\\"


"=\lim_{u\to0}\dfrac{[x(u+\Delta u)2i+y(u+\Delta u)6j]+z(u+\Delta u)3k]-[x(u)i+y(u)j+z(u)k]}{\Delta u}"


"=\lim_{u\to0}\dfrac{ x(u+\Delta u)-x(u)}{\Delta u}2i+\dfrac{ y(u+\Delta u)-y(u)}{\Delta u}6j+\dfrac{ z(u+\Delta u)-z(u)}{\Delta u}3k"


"=\dfrac{dx}{du}2i+\dfrac{dy}{du}6j+\dfrac{dz}{du}3k"